# InstructPipe: Building Visual Programming Pipelines with Human Instructions

### v1

@@ Abstract.
Visual programming provides beginner-level programmers with a coding-free experience to build their customized pipelines. Existing systems require users to build a pipeline entirely from scratch, implying that novice users need to set up and link appropriate nodes all by themselves, starting from a blank workspace. We present InstructPipe, an AI assistant that enables users to start prototyping machine learning (ML) pipelines with text instructions. We designed two LLM modules and a code interpreter to execute our solution. LLM modules generate pseudocode of a target pipeline, and the interpreter renders a pipeline in the node-graph editor for further human-AI collaboration. Technical evaluations reveal that InstructPipe reduces user interactions by 81.1% compared to traditional methods. Our user study (N=16) showed that InstructPipe empowers novice users to streamline their workflow in creating desired ML pipelines, reduce their learning curve, and spark innovative ideas with open-ended commands.
@@
copyright: rightsretainedjournalyear: 2024copyright: rightsretainedconference: Under Review; booktitle: ;doi: XX.XXXX/XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Computing methodologies Visual analyticsccs: Software and its engineering Visual languagesccs: Computing methodologies Machine learning
Figure 1. Existing visual programming requires users to select nodes, conceive pipeline structure, and then create the pipeline starting from an empty workspace. In contrast, InstructPipe facilitates users to build interactive pipelines directly from text instructions.
@@ 1. Introduction
A visual programming interface provides users with a node-graph editor to program. As opposed to writing code in a code editor, the node graph allows users to build a pipeline in a visual workspace with nodes and edges. This approach effectively reduces technical barriers for users to prototype creative applications. Advances in machine learning (ML) further stimulate growing interest in visual programming. Open-sourced ML libraries (e.g., TensorFlow (ten, 2015), PyTorch (Paszke et al., 2019), and Hugging Face (Wolf et al., 2019)) provide users with various encapsulated modules to accelerate AI project development and experimentation. Meanwhile, this also provides valuable protocols for visual programming developers to create systems for ML applications (Du et al., 2023a), where ML practitioners can interactively test “off-the-shelf” models on the node-graph editor. Recent foundational models like large language models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Anil et al., 2023) and findings on Chain-of-Thought (Wei et al., 2023) further stimulate a community-wise interest in visual programming (Wu et al., 2022a, b; Du et al., 2023b), which provides users interactive experiences to explore AI chains.
Despite the development of visual programming platforms in various domains, we observed that existing systems share one similar characteristic: users usually start a creative process in the workspace “from scratch”. This implies that users need to 1) select nodes, 2) ideate the pipeline structure, and finally, 3) connect nodes from a completely empty workspace. For users who are unfamiliar with a particular visual programming platform, such processes can easily overwhelm them, degrading their overall programming experience.
@@
Similar issues also exist when users write programs using text-based editors (there exist many built-in functions in a particular programming language and multiple variables in a program), but advances in LLM assistants show that such challenges can be effectively reduced. For example, GitHub Copilot (GitHub, 2023) makes it possible for users to generate code by simply describing users’ requirements in natural language. Even though the generated code is not absolutely correct, the AI assistance usually finishes a large portion of the task, and programmers may only need to make a few edits to achieve a correct result. To this end, we raise the following questions that motivate our work: How can we build such an assistant to benefit visual programming users?
@@
In this paper, we built InstructPipe, an AI assistant for visual programming users to generate a pipeline through natural language instructions (Figure 1). We implemented InstructPipe on Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines. One major technical challenge in implementing InstructPipe lies in the lack of visual programming data. Different from the standard approach to building a copilot for the text-based editor (i.e., training LLMs using large-scale text-based programs online), it is impractical to collect sufficient data for a particular visual programming platform. We addressed this issue by decomposing the generation process into three steps. Using two separate LLM modules, the system first scopes the potentially useful nodes and then generates pseudocode for a target pipeline. InstructPipe then compiles the pseudocode and renders the pipeline on the node-graph editor to facilitate further user interaction. Our technical evaluation suggests that InstructPipe reduces user interactions by $81.1\%$ compared to building pipelines from scratch. Our system evaluation with 16 participants demonstrated that InstructPipe significantly reduced users’ workload in their creative process. Qualitative results further reveal that InstructPipe effectively supports novices’ on-boarding experience of visual programming systems and allows them to easily prototype a concept for various purposes. As one pioneering work on visual programming copilot, we also observed new challenges caused by humans’ cognitive characteristics and proposed future technical directions toward a next-level, open-ended AI prototyping assistant.
In summary, this work offers the following contributions:
@@
    InstructPipe, an AI assistant that enables users to build ML pipelines from human instructions,
@@
    System design and technical development of InstructPipe, which includes two LLM modules and a code interpreter that generate codes for a visual programming pipeline, compile the code, and render the pipeline in an interactive node-graph editor,
@@
    A technical and a user evaluations that demonstrate effectiveness of InstructPipe, and the corresponding findings that reveal new challenges for the HCI community.
@@ 2.1. Visual Programming
Today, visual programming systems (e.g., LabView (Kodosky, 2020), Unity Graph Editor (Unity, 2023), PromptChainer (Wu et al., 2022a), and Visual Blocks (Du et al., 2023a)) typically feature a node graph editor, providing users with a visual workspace to “write” their program by “building blocks” (Hewett, 2005; Sramek et al., 2023). Recent work further explored the application of visual programming in education (Jiang, 2023; Kovalkov et al., 2020), authoring support (Zhang et al., 2023b; Zhang and Oney, 2020), and robotics (Datta et al., 2012; Huang et al., 2016; Huang and Cakmak, 2017). For example, Zhang et al. (Zhang et al., 2023b) connected the visual programming tool to the concept of teaching by demonstration (Myers, 1986; Li and Landay, 2005; Zhou and Yatani, 2022), allowing users to rapidly customize AR effects in video creation. FlowMatic (Zhang and Oney, 2020) extended traditional visual programming interfaces into 3D virtual environments, providing users with immersive authoring experiences.
More recently, findings on LLM Chains (Wu et al., 2022b) and Chain-of-Thought (Wei et al., 2023) further stimulated researchers to build visual programming tools that chain LLM modules. Developers want to explore various ways to chain an LLM module for various applications. In such a scenario, visual programming provides a great platform for users to focus on high-level exploration. Example research work and industrial products include PromptChainer (Wu et al., 2022a), LangFlow (LangFlow, 2023), and ComfyUI (ComfyUI, 2023).
This work offers technical contributions of implementing and evaluating an AI assistant that supports users to prototype an AI visual programming pipeline from text-based instructions. Compared to a mainstream user workflow in which people begin their creation from scratch, InstructPipe provides a new experience for beginner-level users to easily prototype a pipeline.
2.2. LLMs and Their Applications in Interactive Systems
Early multimodal ML work uses language models to solve Visual Question Answering (VQA) (Andreas et al., 2017; Antol et al., 2015), yet these solutions are limited to simple questions and cannot perform effective reasoning and problem solving (Ma et al., 2023). LLMs revolutionized AI’s reasoning capability (Wei et al., 2023; Zhou et al., 2022) in language, which motivate researchers to build LLM applications in various domains beyond NLP (Surís et al., 2023; Singh et al., 2023; Park et al., 2023). For instance, LLMs augmented the perceptual and planning intelligence in robotics (Singh et al., 2023), supported autonomous driving (Liang et al., 2022), and assisted clinical processes in medical science (Kung et al., 2023; Singhal et al., 2022).
The advances in LLMs empower recent interactive systems (Peng et al., 2023; Park et al., 2022) with enhanced machine intelligence. Recent research leverages LLMs to edit visualization (Wang et al., 2023a; Shen et al., 2023), receive AI explanation (Wu et al., 2021), facilitate communications (Liu et al., 2023), understand user interface (Wang et al., 2023b), and study simulated social behaviors (Park et al., 2023). The revolution also motivated HCI researchers to design new interfaces for LLMs (Jiang et al., 2023; Suh et al., 2023). Graphologue (Jiang et al., 2023) augmented LLM responses by displaying interactive diagrams along with the response text, which depicts the semantic logic in a paragraph. Sensecape (Suh et al., 2023) provides users with a workspace to explore long LLM responses in a hierarchical structure.
InstructPipe focuses on utilizing LLMs for generating pipelines in visual programming with human instructions (Hu et al., 2023; Surís et al., 2023). This work shares a similar vision with Prompt2Model (Viswanathan et al., 2023) and VisProg (Gupta and Kembhavi, 2023). Prompt2Model (Viswanathan et al., 2023) finetunes a BERT model (Devlin et al., 2019) using data generated by instructions. VisProg (Gupta and Kembhavi, 2023) produces Python code from instructions with task-dependent few-shot prompting. However, prior arts focus on full AI automation and lack an interactive workspace for human-AI collaboration. In contrast, InstructPipe generates and compiles a pipeline (without finetuning), while rendering the pipeline in a visual programming interface, facilitating an interactive, collaborative, and explainable workflow.
@@ 3.1. User Workflow
To generate a pipeline, users can first click the “InstructPipe” button on the top-right corner of the interface (1(b)). The system then activates a simple dialog (1(a)) in which users can 1) provide a description of their target pipeline and 2) tag the pipeline. The tag can be “language”, “visual”, or “multimodal”. After users click the “Submit” button in the dialog, InstructPipe renders a visual programming pipeline on the node-graph editor. Based on the result, users can further refine the pipeline in the visual programming interface.
@@ 3.2. Primitive Nodes
Figure 3. An overview of major primitive processor nodes supported by InstructPipe. 17 of the 20 processor nodes are categorized by the type of accepted I/O into a $3\times 3$ matrix. Note that “PaLM” represents two nodes in InstructPipe, i.e., a text generation model and a chat model of PaLM (Anil et al., 2023).
InstructPipe supports 27 primitive nodes in Visual Blocks, including seven I/O nodes (e.g., live camera and markdown viewer) and 20 processor nodes. Appendix A provides a description for each of the 27 nodes. Compared to related work (Gupta and Kembhavi, 2023; Surís et al., 2023) that automates ad hoc ML inferences in specific use scenarios, we aim at an open-ended use case with diverse primitive nodes. Figure 3 visualizes 17 (out of 20 in total) processor nodes according to the I/Os of the nodes. The remaining three nodes are “Google Sheet” (which takes a Google Sheet URL as input and outputs the sheet data), “image_mixer” (which combines multiple images), and “virtual_sticker” (which casts a sticker on a person’s face on a live camera). “Metadata” in Figure 3 indicates intermediate data used in ML pipelines, which can be a segmentation mask that describes an input image, or a URL that describes a target news that users want to read. As shown in the matrix, InstructPipe contains a wide range of nodes that supports the creation of complex ML pipelines.
@@ 4. Pipeline Generation from Instructions
InstructPipe leverages LLMs to generate visual programming pipelines from text instructions. There exist two prevailing approaches for customizing LLMs: 1) fine-tuning LLMs (Brooks et al., 2023; Liu et al., 2023), and 2) few-shot prompting (Park et al., 2023; Dong et al., 2023). For our task, fine-tuning LLMs requires a substantial volume of annotated data comprising pairs of pipelines and prompts. Additionally, a growing list of nodes poses new challenges to scaling this approach with new data annotations. In comparison, few-shot prompting can seem more practical (Gupta and Kembhavi, 2023; Wei et al., 2023; Yao et al., 2023), but it is challenging to design an efficient prompt that fits within a reasonable number of tokens. The configuration file of the 27 nodes alone includes 8.2k tokens. Moreover, because of the combinatorial explosion of the 27 nodes in the system, it is not clear how many prompt examples are needed and how we can construct in-context pipeline examples.
@@
Figure 4 illustrates the high-level workflow of the InstructPipe implementation. InstructPipe includes two LLM modules (highlighted in red): 1) a Node Selector (section 4.2) and 2) a Code Writer (section 4.3). Given a user instruction and a pipeline tag, we first devise the Node Selector to identify a list of potential nodes that would be used according to the instruction. In the Node Selector, we prompt the LLM with a very brief description of each node, aiming to filter out unrelated nodes for a target pipeline. The selected nodes and the original user input (the prompt and the tag) are then fed into the Code Writer, which generates pseudocode for the desired pipeline. In Code Writer, we provide the LLM with detailed description and examples of each selected nodes to ensure LLMs have a thorough understanding of each candidate node. Finally, we employ a Code Interpreter to parse the pseudocode and render a visual programming pipeline for the user to interact with.

### v2

@@ Abstract.
Visual programming has the potential of providing novice programmers with a low-code experience to build customized processing pipelines. Existing systems typically require users to build pipelines from scratch, implying that novice users are expected to set up and link appropriate nodes from a blank workspace. In this paper, we introduce InstructPipe, an AI assistant for prototyping machine learning (ML) pipelines with text instructions. We contribute two large language model (LLM) modules and a code interpreter as part of our framework. The LLM modules generate pseudocode for a target pipeline, and the interpreter renders the pipeline in the node-graph editor for further human-AI collaboration. Both technical and user evaluation (N=16) shows that InstructPipe empowers users to streamline their ML pipeline workflow, reduce their learning curve, and leverage open-ended commands to spark innovative ideas.
@@
copyright: rightsretainedjournalyear: 2024copyright: rightsretaineddoi: XX.XXXX/XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Computing methodologies Visual analyticsccs: Software and its engineering Visual languagesccs: Computing methodologies Machine learning
@@ 1. Introduction
A visual programming interface provides users with a node-graph editor to program through interaction with visual elements. As opposed to writing code in a code editor, the node graph allows users to design pipelines by configuring nodes and connecting them with edges in a visual workspace. This alternative user interface approach often accelerates experimentation and exploration in the prototyping phases of creative applications, and can make advanced technology more accessible to beginners. Advances in machine learning (ML) further stimulate growing interest in visual programming. Open-source ML hubs (e.g., TF-Hub (ten, 2015), PyTorch-Hub (Paszke et al., 2019), and Hugging Face (Wolf et al., 2019)) contribute large numbers of encapsulated modules that accelerate AI project development and experimentation, and such libraries provide important resources for an ML-based visual programming platform. Recent advancements in large language models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Anil et al., 2023) and findings on Chain-of-Thought (Wei et al., 2023) have further stimulated community-wide interest in visual programming (Wu et al., 2022a, b; Du et al., 2023b; Arawjo et al., 2024), suggesting further potential in the interactive exploration of AI chains.
@@
Despite the development of visual programming platforms in various domains, we observed that existing systems share one similar characteristic: users usually initiate a creative process in the workspace “from scratch”. This implies that users need to 1) select nodes, 2) ideate the pipeline structure, and finally, 3) connect nodes within a completely empty workspace. For users unfamiliar with a particular visual programming platform, such processes can be overwhelming, degrading their overall programming experience. Providing pipeline templates may reduce on-boarding efforts (Chen et al., 2024; Fogarty et al., 2001), but the templates inherently lack flexibility and are not easily adaptable to users’ specific needs. Similar issues also arise when users write programs using text-based editors (there exist many built-in functions in a particular programming language and multiple variables in a program), but advances in LLM assistants show that such challenges can be effectively reduced. For example, GitHub Copilot (GitHub, 2023) enables users to generate code by simply describing users’ requirements in natural language. Even though the generated code is not absolutely correct, the AI assistance usually finishes a large portion of the task, and programmers may only need to make a few edits to achieve a correct result (Chris Perry, 2023; Jupyter-ai, 2024). To this end, we raise the following question that motivates our work: How can we build visual programming assistants to accelerate design and prototyping of ML pipelines?
@@
This paper introduces InstructPipe, a visual programming AI assistant that enables ML pipeline generation and design through natural language instructions. InstructPipe facilitates node connection and selection, allowing users to focus on more creative tasks like parameter tuning and interactive analysis within the visual programming workspace. We focus our AI assistant exploration on ML-based pipelines, and therefore implemented InstructPipe as an extension to Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines. One major technical challenge in implementing InstructPipe lies in the lack of visual programming data, making it impractical to finetune a dedicated code-LLM similar to how developers build text-editor-based copilots (GitHub, 2023; Chris Perry, 2023; Jupyter-ai, 2024). We addressed this issue by decomposing the generation process into three steps (Fig. 1). InstructPipe’s first LLM module scopes the potentially useful nodes, while the second LLM module generates pseudocode for a pipeline. InstructPipe then parses the pseudocode and renders the pipeline in the node-graph editor to facilitate further user interaction. Our technical evaluation suggests that InstructPipe reduces the necessary user interactions by $81.1\%$ when users select and connect nodes, compared to building them from scratch. This can potentially streamline the development process, and allows users to focus on more novice-friendly interactions like parameter-tuning and human-in-the-loop verification. Our system evaluation with 16 participants demonstrated that InstructPipe significantly reduced users’ workload in their creative process. Qualitative results further reveal that InstructPipe effectively supports novices’ on-boarding experience of visual programming systems and allows them to easily prototype concepts for various purposes. In our experiments, we also observed new challenges caused by human cognitive characteristics, and propose future technical directions towards open-ended AI prototyping assistants.
In summary, we contribute:
@@
    InstructPipe, a visual programming AI assistant that enables users to generate ML pipelines from human instructions by automating node selection and connection,
@@
    System design and technical development of InstructPipe. The system consists of two LLM modules and a code interpreter, which generate the specification for the visual programming pipeline, compile the code, and render the pipeline in an interactive node-graph editor,
@@
    Technical and user evaluations that characterize the effectiveness of InstructPipe, and contribute findings that reveal new challenges for the HCI community.
@@ 2.1. Visual Programming
Today, visual programming systems (e.g., LabView (Kodosky, 2020), Unity Graph Editor (Unity, 2023), PromptChainer (Wu et al., 2022a), ComfyUI (ComfyUI, 2023) and Visual Blocks (Du et al., 2023a)) typically feature a node graph editor, providing users with a visual workspace to “write” their program using “building blocks” (Hewett, 2005; Sramek et al., 2023; Yu et al., 2024). Recent work further explored the application of visual programming in education (Jiang, 2023; Kovalkov et al., 2020; Chen et al., 2024), XR creativity support (Zhang et al., 2023b; Zhang and Oney, 2020; Ye et al., 2024), and robotics (Datta et al., 2012; Huang et al., 2016; Huang and Cakmak, 2017). For example, Zhang et al. (Zhang et al., 2023b) connected the visual programming tool to the concept of teaching by demonstration (Myers, 1986; Li and Landay, 2005; Zhou and Yatani, 2022), allowing users to rapidly customize AR effects in video creation. FlowMatic (Zhang and Oney, 2020) extended traditional visual programming interfaces into 3D virtual environments, providing users with immersive authoring experiences.
Advancements in AI has introduced many repositories of advanced ML models (Shen et al., 2024; HuggingFace, 2022), and an increasing number of researchers are exploring AI chains (Wu et al., 2022b; LangChain, 2023). This progress has motivated HCI researchers to design and build a range of visual programming interfaces to support the AI development process (Wu et al., 2022a; LangFlow, 2023; ComfyUI, 2023). For example, ChainForge is a web-based platform for developers to explore various LLM-related configuration and designs in a wide range of applications (Arawjo et al., 2024). Visual Blocks enables creation and interaction of advanced ML pipelines that can leverage state-of-the-art computer vision and computer graphics models in the browser (Du et al., 2023a).
This work contributes the technical system, implementation and evaluation of a novel AI assistant that enables the use of text-based instructions in visual programming of ML pipelines. Compared to typical workflows in which people manually build their pipelines, InstructPipe has the potential to accelerate ML pipeline prototyping in visual programming.
2.2. Interactive Systems with LLMs
The advances in LLMs bring many research directions for HCI researchers. Researchers have started designing new LLM interfaces, to advance beyond the currently dominant chatbot interface (e.g., OpenAI ChatGPT, Google Gemini). For example, Graphologue (Jiang et al., 2023) augmented LLM responses with interactive diagrams that visualize response texts in a structured format. Sensecape (Suh et al., 2023) provides users with a workspace to explore long LLM responses in a hierarchical structure.
Many HCI researchers integrated LLMs in conventional interactive systems and demonstrated that such enhanced machine intelligence can provide new user experiences (Peng et al., 2023; Park et al., 2022; Liu et al., 2023a; Wang et al., 2023b). This research principle is widely applied in many downstream HCI applications including visualization (Wang et al., 2023a; Shen et al., 2023), explainable AI (Wu et al., 2021; Wang et al., 2024), and social science (Park et al., 2023; Liu et al., 2023b). For example, Chen et al. (Chen et al., 2023) utilized LLMs to bridge low-level sensor information with high-level human requests. Experiments showed that such connection allows users to “construct their personalized contexts [for an intelligent system] more quickly, accurately, and naturally”. To interface human intention with machine operations, researchers typically utilized LLMs by following the ReAct (reasoning and acting) paradigm (Yao et al., 2023). For example, Park et al. (Park et al., 2023) simulated human behaviors in an artifical social system by leveraging LLMs as intelligent agents that can perceiving the environment, planning their behaviors, and acting in the environment. Automated Visualization (AutoViz) researchers employ LLMs for data analysis and reasoning for presenting the visualization (Narechania et al., 2021; Mitra et al., 2022; Sah et al., 2024). For example, LIDA features four modules in the visualization pipeline to 1) summarize a structured dataset, 2) explore user’s goal, 3) generate code for visualization, and 4) render visualization (Dibia, 2023). ChartGPT further constructs a dedicated dataset for chart visualization, and finetunes an LLM for fully automating the data visualization pipeline (Tian et al., 2024).
InstructPipe extends the application of ReAct-like LLM frameworks (Tian et al., 2024; Dibia, 2023) to visual programming and demonstrates its effectiveness to support rapid prototyping with lower user workload. Additionally, introducing visual programming to the ReAct framework showcases an interface solution for human-AI collaboration. That being said, our work values partially correct AI generation, though the previous literature considers it as a complete generation failure (Dibia, 2023; Gupta and Kembhavi, 2023). We render the imperfect results in a visual programming platform, so that even novices can intuitively stand on the shoulder of the AI generation for their work.
@@ 3.1. User Workflow
To generate a pipeline, users first click the “InstructPipe” button in the top-right corner of the interface (2(b)). The system then activates a simple dialog (2(a)) in which users provide a description and a tag for their desired pipeline. The tag can be “language”, “visual”, or “multimodal”, and helps guide the pipeline generation. After users click the “Submit” button, InstructPipe generates a visual pipeline in the node-graph editor. More specifically, InstructPipe generates a directed acyclic graph (DAG) of a visual programming pipeline. This implies that it uses default node parameters (e.g., the “temporature” or “max_tokens” value of an LLM node). Therefore, after the generation, the user needs to 1) finish the graphic structure if necessary, and 2) perform parameter tuning as well as human-in-the-loop evaluation of the pipeline quality interactively in the visual programming platform. As we will show in our evaluation, this new human-AI collaboration approach reduces users’ workload on the technical portion of the visual programming tasks (selecting and connecting nodes), and thus provide a more novice-friendly experiences for technical visual programming platforms.
@@ 3.2. Primitive Nodes
Figure 3. The distribution of 20 primitive processor nodes supported by InstructPipe. Note that “PaLM” represents two nodes in InstructPipe, i.e., a text generation model and a chat model of PaLM (Anil et al., 2023).
InstructPipe supports 27 primitive nodes in Visual Blocks. We achieved this node library of InstructPipe by filtering out nodes without explicit definition of their functions [^1] Note that Visual Blocks is a system that is actively being updated, and there are more nodes now.. For example, ‘TFLite model runner’ is an implicitly defined node: the user needs to input a tensorflow hub link to define its functionality. As we mentioned previously, InstructPipe focuses on generating a DAG and leaves the parameter tuning task to users. Adding such implicit nodes without clear definition of the functionality can easily confuse our AI assistant in the generation process, and thus we decide to exclude these nodes in the node library of InstructPipe.
The 27 nodes in our library include three input nodes, four output nodes and 20 processor nodes. The following shows an example node in each category, and we leave the full node library description in Appendix A:
  • “live camera” (an input node): Capture video stream through your device camera
  • “markdown viewer” (an output node): Render Markdown strings into stylized HTML.
  • “imagen” (a processor node): Generate an image based on a text prompt.
We distributed 20 processor nodes based on the data type of its I/O edges and visualized it in Fig. 3. For example. “Google Web Search” takes “Texts” information as input and output new “Texts”, and “OCR” takes an image (vision-based information) as input and output “Texts”. “Features” in Fig. 3 indicates a wide range of intermediate data format used in ML pipelines, including segmentation masks, a pose landmarks, URLs and etc. As shown in the matrix, InstructPipe contains a wide range of nodes that supports the creation of complex ML pipelines. Compared to related work that automates ad hoc ML inferences in specific use scenarios (Gupta and Kembhavi, 2023; Surís et al., 2023), InstructPipe makes one more step towards the open-ended assistants with a more diverse set of primitive nodes. Further extending our node library can effectively empower the capability of our AI assistant, which we leave as critical future work. In the current implementation of InstructPipe, we focus on demonstrating its capability based on our focused node library, and explore what new experiences this AI assistant can bring to our users.
@@ 4. Pipeline Generation from Instructions
InstructPipe leverages LLMs to generate visual programming pipelines from text instructions. There are two prevailing approaches for LLM-customization, fine-tuning (Qin et al., 2023b; Liu et al., 2023a), and few-shot prompting (Park et al., 2023; Gupta and Kembhavi, 2023). Fine-tuning would require a substantial volume of annotated data, with pairs of pipelines and prompts, and it is hard to achieve for a specific visual programming platform. Additionally, a growing list of nodes would consistently require 1) new data annotation and 2) retraining the model, making this approach less sustainable. In comparison, few-shot prompting is a more practical approach for prototyping an interaction concept to understand the new experience it would bring to the community (Gupta and Kembhavi, 2023; Wei et al., 2023; Yao et al., 2023). One major challenge of applying LLMs in visual programming AI assistant lies in designing efficient prompts that fit within a reasonable number of tokens. Even though we focus our exploration on 27 nodes, the node configuration file alone includes 8200 tokens. Further formulating pipeline examples as in-context few-shot examples would result in a combinatorial explosion, causing an overwhelming number of tokens in the prompt.
@@
Fig. 1 illustrates the high-level workflow of the InstructPipe implementation. InstructPipe leverages two LLM modules (highlighted in red); a Node Selector (section 4.2), and a Code Writer (section 4.3). Given a user instruction and a pipeline tag, we first devise the Node Selector to identify a list of potential nodes that would be used according to the instruction. In the Node Selector, we prompt the LLM with a very brief description of each node, aiming to filter out unrelated nodes for a target pipeline. The selected nodes and the original user input (the prompt and the tag) are then fed into the Code Writer, which generates pseudocode for the desired pipeline. In Code Writer, we provide the LLM with detailed description and examples of each selected nodes to ensure the LLM has extensive context for each candidate node. Finally, we employ a Code Interpreter to parse the pseudocode and render a visual programming pipeline for the user to interact with.